OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 2x - 3y - z = -1 x - 4y + 3z = -11 2x + 9y + 3z = -13 -11y + 5z = 21 y + 9z = 9 5y - 7z = 21 17y - 3z = 9 7x - 13y = -14 8x + 0y = -16 -5x +5y = -8 -4x + 18z = -10 I don't understand how to get these at all..

OpenStudy (mathstudent55):

In the first two choices, there is no x variable. That means they solved one equation for x and substituted that value of x into the other two. Look at the three equations, the second can be easily solved for x. x - 4y + 3z = -11 x = 4y - 3z - 11 Now substitute this value of x into the first and third equations. First equation: 2x - 3y - z = -1 2(4y - 3z - 11) - 3y - z = -1 8y - 6z - 22 - 3y - z = -1 5y - 7z = 21 Notice the above equation is included in choice B. Third equation: 2x + 9y + 3z = -13 2(4y - 3z - 11) + 9y + 3z = -13 8y - 6z - 22 + 9y + 3z = -13 17y -3z = 9 This is the other equation of choice B.

OpenStudy (anonymous):

thanks you stated it so clearly! thanks so much

OpenStudy (mathstudent55):

you're welcome